After declining between 1940 and 1980, the number of multigenerational households has been increasing since 1980. The function ho) 0012x2-0.573x+35 727 can be used to estimate the number of multigenerational households in country A, in millions, x years after 1940 In what year were there 50 million multigenerational households? There were 50 million multigenerational households in (Round down to the nearest whole number.)
Answer #1

\\h(x) = 0.012x^2-0.573x+35.727\\ \\to\ cal\ the\ year\ in\ which\ hh\ were\ 50\ million\ put\ h(x) = 50\\ \\h(x) = 0.012x^2-0.573x+35.727 = 50\\ \\h(x) = 0.012x^2-0.573x+14.273 = 0\\

\\x = \frac{0.573 \pm \sqrt{0.573^2 - 4*14.273*0.012}}{2*0.012}\\ \\x = -18.0706\ and\ 65.8206\\ \\it\ means\ in\ year\ 2005\ the\ hh\ were\ 50\ million\ but\ in\ that\\ \\year\ the\ equation\ is\ not\ valid\\

so answer is in none of the years.